40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Nov 26 19:38:40 2017, Steve, Update, VAC, TP3 drypump replaced again after_replacement.png
    Reply  Thu Nov 30 10:38:39 2017, Steve, Update, VAC, annuloses are not pumped AnnulosesNotPumped.png
Message ID: 13454     Entry time: Sun Nov 26 19:38:40 2017     In reply to: 0     Reply to this: 13459
Author: Steve 
Type: Update 
Category: VAC 
Subject: TP3 drypump replaced again 

The TP3 foreline pressure was 4.8 Torr, 50K rpm 0.54A and 31C........Maglev rotation normal 560 Hz.......    IFO pressure 7.2e- 6 Torrit was not effected

V1 closed ......replaced drypump.........V1 opened

IFO 6.9e-6 Torrit at 19:55, TP3fl 18 mT,  50Krpm 0.15A 24C

VM1 is still closed

 

Attachment 1: after_replacement.png  24 kB  | Hide | Hide all
after_replacement.png
ELOG V3.1.3-